# Utility-Diversity Aware Online Batch Selection for LLM Supervised Fine-tuning

### v2

@@ 1 Introduction
The rapid progress of Large Language Models (LLMs) has reshaped natural language processing and enabled impressive generalization across a wide range of domains (achiam2023gpt; brown2020language; liu2024deepseek; touvron2023llama; bai2023qwen). To further improve their performance in specialized areas, Supervised Fine-tuning (SFT) has emerged as a typical post-training paradigm. However, training on a full dataset is not always advantageous. It entails substantial computational cost and has been shown to be less effective than using a small amount of carefully curated data (albalak2024survey; xia2024less; zhou2023lima). These challenges highlight the importance of principled data selection strategies that can improve both efficiency and effectiveness in SFT.
In this work, we address this challenge by focusing on online batch selection, a paradigm that dynamically evaluates sample value and performs filtering during training (wang2024greats; loshchilov2015online; katharopoulos2018not; mindermann2022prioritized). Instead of using all samples in a batch, the model assesses the importance of each sample as it is encountered and selects only a subset to participate in parameter updates. This approach enables the selection process to adapt in real time to the model’s current state and learning trajectory, improving both training efficiency and effectiveness.
However, despite promising progress, current online batch selection methods still face several issues that hinder their practical deployment. These methods typically rely solely on data utility, e.g., picking sample subsets with high loss (loshchilov2015online; jiang2019accelerating) or gradient magnitude (katharopoulos2018not). While such heuristics enable capturing the critical samples to reduce the current training loss, they evaluate sample value from a limited perspective. In practice, effective selection also requires considerations of intra-sample diversity, to reflect the richness of information with fewer repetitive phrases within each training instance, and inter-sample diversity to suppress redundancy across examples by avoiding repeated training on near-duplicate content (lee2021deduplicating; tirumala2023d4). Popular literature (loshchilov2015online; katharopoulos2018not; mindermann2022prioritized) prioritizes data utility and rarely examines the selection criteria through the lens of diversity.
Another primary issue involves both external dependencies and computational overhead. Some approaches (mindermann2022prioritized; wang2024greats) rely on a held-out validation set, which may be unavailable since we hardly know the exact distribution of test dataset. Other methods (mindermann2022prioritized; deng2023towards) use a reference model that is also shown to be impractical in the real world (kaddour2023no). Besides these external dependencies, several selection algorithms (katharopoulos2018not; wang2024greats; mindermann2022prioritized) introduce substantial computational costs that even exceed the expense of full-dataset training, raising efficiency concerns.
@@
Leveraging logits as basis for online scoring. To operationalize these desiderata, a key challenge lies in how to efficiently and reliably score candidate samples during training. In online batch selection, each candidate sample must be evaluated under the current model to quantify its contribution to learning. This evaluation typically requires at least a forward pass, as it can fully capture how the model presently interprets the sample and thus provide an accurate basis for selection (huang2023look; shorinwa2025survey). To achieve this efficiently (D3) without external resources (D2), we must avoid expensive gradient computations for each candidate sample and leverage intrinsic signals already available during forward propagation. These considerations collectively justify leveraging the model’s output logits, which are naturally produced during the forward pass and encode rich information about both sample utility and diversity (qiu2024semantic; geng2023survey). Specifically, we compute an intra-sample importance score based on the nuclear norm of the logits, which captures the utility and intra-sample diversity of a sample, and an inter-sample importance score using low-dimensional similarity matching against previously selected samples to quantify inter-sample diversity (D1). Integrating these scores allows the training process to balance exploitation of high-utility samples with exploration of less-visited regions in the data distribution, thereby improving the overall effectiveness of online batch selection.
@@
    To capture intra-sample value, we leverage the nuclear norm of the logits matrix, which naturally reflects both data utility and intra-sample diversity, providing a principled criterion for assessing within-sample informativeness.
@@
    To estimate inter-sample diversity, we design a structured bilinear random projection of logits for compact embedding and show that similarity matching with historical data points effectively reduces redundancy with negligible overhead.
@@
    Our UDS operates directly on forward-pass outputs without external resources, achieving faster convergence than full-dataset SFT and surpassing existing online batch selection baselines.
@@ 2.1 Problem Formulation of Online Batch Selection
At training iteration $t$, we draw a candidate batch $\mathcal{B}_{t}=\{(\bm{x}_{t}^{i},y_{t}^{i})\}_{i=1}^{B}$ from the corpus. The goal of online batch selection is to choose a subset $\widehat{\mathcal{B}}_{t}\subseteq\mathcal{B}_{t}$ that maximizes the immediate (or near-term) benefit to the model. Let $s(\bm{x}_{t}^{i};\bm{\theta}_{t},\mathcal{H}_{t})\in\mathbb{R}$ denote a per-sample importance score computed under the current model state $\bm{\theta}_{t}$ and optional history $\mathcal{H}_{t}$ (e.g., a small buffer of recent embeddings). The selection objective can be viewed as an optimization problem:
@@
where $K$ controls the fraction of data we aim to select.
@@ 2.2 Autoregressive Language Modeling
During fine-tuning, the parameters $\bm{\theta}_{t}$ are updated to maximize the likelihood of training sequences. At iteration $t$, the online selection mechanism identifies an informative subset $\widehat{\mathcal{B}}_{t}$ from the candidate batch $\mathcal{B}_{t}$, which is then used to perform a gradient update $\bm{\theta}_{t}\to\bm{\theta}_{t+1}$.
@@ 3 Utility-Diversity Sampling (UDS)
We propose UDS, an online batch selection framework that jointly considers data utility, intra-sample diversity, and inter-sample diversity. It scores and samples each data point using a mixture of two complementary scores: (i) the Nuclear Norm of the logits matrix ($s_{\text{intra}}^{t,i}$), capturing optimization utility and intra-sample diversity; and (ii) the Diversity Distance ($s_{\text{inter}}^{t,i}$), measuring dispersion against recent selections. An overview is shown in Figure 1, with pseudocode provided in Algorithm 1.
@@ 3.1 Intra-sample Importance Score via Nuclear Norm
The intra-sample importance score ($s_{\text{intra}}^{t,i}$) can be characterized through two complementary views: (i) optimization utility—how much the sample can contribute to loss reduction during training, and (ii) intra-sample diversity—how diverse the model’s token-level outputs are within the sequence. We employ the nuclear norm (trace norm) of the logits matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ to capture both aspects. The nuclear norm $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ is defined as the sum of singular values of $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$. Let $\{\sigma_{1},\sigma_{2},\dots,\sigma_{r}\}$ denote its singular values (with $r=\mathrm{rank}\left(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\right)$), then $s_{\text{intra}}^{t,i}$ is computed as:
@@
In mathematics, the nuclear norm and the Frobenius norm can bound each other, as shown in Lemma 3.1. A detailed explanation is provided in Appendix C.1.
@@
(horn2012matrix) For any matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$, the following inequality holds:
$\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}\leq\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}\leq\sqrt{\min(N,V)}\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm, computed as the square root of the sum of squares of all entries in the matrix. The left inequality achieves equality when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-$1$ and has only one non-zero singular value. The right inequality achieves equality when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank and all singular values are equal.
Thus, a larger nuclear norm arises in two ways: (1) the Frobenius norm increases, indicating larger logits, and (2) for a fixed Frobenius norm, the nuclear norm shifts closer to the upper bound. Below we clarify how these two effects embody the two complementary aspects of sample-importance score.
@@
1:Candidate batch $\mathcal{B}_{t}$ ($t=0,1,\ldots,T$), initial model $\pi_{\bm{\theta}_{0}}$, memory queue $\bm{Q}$
2:Initialize memory queue $\bm{Q}\leftarrow\{\}$
3:for $t=0,1,\ldots,T$ do
4:  // Calculate importance score and dynamically sampling
5:  for each sample $\bm{x}_{t}^{i}\in\mathcal{B}_{t}$ do
6:   Perform forward pass to obtain logits $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\leftarrow\pi_{\bm{\theta}_{t}}(\bm{x}_{t}^{i})$
7:   Calculate intra-sample importance score $s_{\text{intra}}^{t,i}\leftarrow\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$
8:   Randomly project to low-dimensional embeddings $\bm{z}_{t}^{i}\leftarrow\operatorname{vec}(\bm{\Gamma}_{2}\cdot\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\cdot\bm{\Gamma}_{1}^{\top})$
9:   Calculate inter-sample importance score $s_{\text{inter}}^{t,i}\leftarrow\frac{1}{|\bm{Q}|}\sum_{\bm{z}_{j}\in\bm{Q}}\|\bm{z}_{t}^{i}-\bm{z}_{j}\|_{2}$
10:   Combine intra- and inter-sample importance scores $s_{\text{total}}^{t,i}=s_{\text{intra}}^{t,i}+\alpha*s_{\text{inter}}^{t,i}$
11:  end for
12:  Dynamically select top-$K$ samples $\widehat{\mathcal{B}}_{t}\leftarrow\operatorname{TopK}(\{s_{\text{total}}^{t,i}\}_{i=1}^{B},K)$
13:  // Update memory queue
14:  while $|\bm{Q}|+K>M$ do
15:   $\bm{Q}\leftarrow\bm{Q}\setminus\{\text{oldest element}\}$
16:  end while
17:  $\bm{Q}\leftarrow\bm{Q}\cup\{\bm{z}_{t}^{i}\mid\bm{x}_{t}^{i}\in\widehat{\mathcal{B}}_{t}\}$
18:  Update model parameters $\bm{\theta}_{t+1}\leftarrow\bm{\theta}_{t}$ through backpropagation
19:end for
Larger Nuclear Norm Positively Correlates With Higher Optimization Utility.
Traditional notions of sample difficulty—such as maximum loss (loshchilov2015online) or maximum gradient (katharopoulos2018not)—often misalign with actual training dynamics (wang2024greats). Instead, we characterize sample importance through its potential contribution to loss reduction, termed its optimization utility. For ease of exposition, we assume training with SGD using batch size $B$ and a learning rate $\eta_{t}$. An extension to Adam is provided in Appendix C.2. The parameter update at iteration $t$ is $\bm{\theta}_{t+1}-\bm{\theta}_{t}=-\frac{\eta_{t}}{B}\sum_{j=1}^{B}\nabla_{\bm{\theta}}\ell(\bm{x}_{t}^{j};\bm{\theta}_{t})$. After the update, the logits matrix for datapoint $\bm{x}_{t}^{i}$ can be expanded using a first-order Taylor expansion:
@@
$\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})=\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t+1})-\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\approx-\frac{\eta_{t}}{B}\nabla_{\bm{\theta}}\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\cdot\left(\sum_{j=1}^{B}\nabla_{\bm{\theta}}\ell(\bm{x}_{t}^{j};\bm{\theta}_{t})\right).$ (6)
Rather than reasoning in parameter space, we can view the training step as producing a logits perturbation $\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$. The first-order Taylor expansion of the loss around $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ gives
$\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t}):=\ell\left(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})+\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\right)-\ell\left(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\right)\approx\langle\nabla_{\bm{L}}\ell(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})),\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\rangle,$ (7)
@@
Figure 2: Strong correlation between $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$ and $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ during the training process. Each point represents the correlation coefficient calculated from $B$ pairs of values within a single batch. We extract the result using Qwen-2.5-7B on MMLU.
Figure 3: Strong correlation between $rank(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t}))$ and $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ during the training process. Each point represents the correlation coefficient calculated from $B$ pairs of values within a single batch. We extract the result using Qwen-2.5-7B on MMLU.
Intuitively, as the Frobenius norm of the logits $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$ increases, the perturbation $\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ also tends to grow uniformly. This occurs because each entry in the activation logits becomes larger, and these values are directly used during backpropagation. In contrast, $\bm{\Delta}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is less sensitive to scale because it depends only on the normalized probabilities and the one-hot labels. Therefore, $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$, $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$, and the attainable loss reduction $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$ typically grow in tandem. Empirically, Figure 2 confirms a strong correlation between a sample’s loss reduction and its nuclear norm. Hence, the nuclear norm can serve as an indicator of a sample’s optimization utility.
When a single sample induces a larger loss reduction, it indicates that the sample both challenges the model’s current predictions and provides informative gradients that effectively guide parameter updates. Such samples are particularly valuable for data selection: they not only highlight under-learned regions of the input space, but also accelerate training by yielding stronger optimization signals compared to redundant or already well-learned samples. We provide a categorization of the loss before and after training in Table 2. Hence, we prioritize selecting samples with larger nuclear norm that are more likely to have higher loss reduction.
@@
High Too Hard Informative

### v3

@@ 1 Introduction
The rapid progress of Large Language Models (LLMs) has reshaped natural language processing and enabled impressive generalization across a wide range of domains (Achiam et al., 2023; Brown et al., 2020; Liu et al., 2024; Touvron et al., 2023; Bai et al., 2023). To further improve their performance in specialized areas, Supervised Fine-tuning (SFT) has emerged as a typical post-training paradigm. However, training on a full dataset is not always advantageous. It entails substantial computational cost and has been shown to be less effective than using a small amount of carefully curated data (Albalak et al., 2024; Xia et al., 2024; Zhou et al., 2023). These challenges highlight the importance of principled data selection strategies that can improve both efficiency and effectiveness in SFT.
In this work, we address this challenge by focusing on online batch selection, a paradigm that dynamically evaluates sample value and performs filtering during training (Wang et al., 2024; Loshchilov and Hutter, 2015; Katharopoulos and Fleuret, 2018; Mindermann et al., 2022). Instead of using all samples in a batch, the model assesses the importance of each sample as it is encountered and selects only a subset to participate in parameter updates. This approach enables the selection process to adapt in real time to the model’s current state and learning trajectory, improving both training efficiency and effectiveness.
However, despite promising progress, current online batch selection methods still face several issues that hinder their practical deployment. These methods typically rely solely on data utility, e.g., picking sample subsets with high loss (Loshchilov and Hutter, 2015; Jiang et al., 2019) or gradient magnitude (Katharopoulos and Fleuret, 2018). While such heuristics enable capturing the critical samples to reduce the current training loss, they evaluate sample value from a limited perspective. In practice, effective selection also requires considerations of intra-sample diversity, to encourage token-level exploration with fewer repetitive phrases within each training instance (Xie et al., 2020), and inter-sample diversity to suppress redundancy across examples by avoiding repeated training on near-duplicate content (Lee et al., 2021; Tirumala et al., 2023). Popular literature (Loshchilov and Hutter, 2015; Katharopoulos and Fleuret, 2018; Mindermann et al., 2022) prioritizes data utility and rarely examines the selection criteria through the lens of diversity.
@@
Another primary issue involves both external dependencies and computational overhead. Some approaches (Mindermann et al., 2022; Wang et al., 2024) rely on a held-out validation set, which may be unavailable since we hardly know the exact distribution of test dataset. Other methods (Mindermann et al., 2022; Deng et al., 2023) use a reference model that is also shown to be impractical in the real world (Kaddour et al., 2023). Besides these external dependencies, several selection algorithms (Katharopoulos and Fleuret, 2018; Wang et al., 2024; Mindermann et al., 2022) introduce substantial computational costs that even exceed the expense of full-dataset training, raising efficiency concerns.
Leveraging logits as basis for online scoring. To operationalize these desiderata, a key challenge lies in how to efficiently and reliably score candidate samples during training. In online batch selection, each candidate sample must be evaluated under the current model to quantify its contribution to learning. This evaluation typically requires at least a forward pass, as it can fully capture how the model presently interprets the sample and thus provide an accurate basis for selection (Huang et al., 2023; Shorinwa et al., 2025). To achieve this efficiently (D3) without external resources (D2), we must avoid expensive gradient computations for each candidate sample and leverage intrinsic signals already available during forward propagation. These considerations collectively justify leveraging the model’s output logits, which are naturally produced during the forward pass and encode rich information about both sample utility and diversity (Qiu and Miikkulainen, 2024; Geng et al., 2023). Specifically, we compute an intra-sample importance score based on the nuclear norm of the logits, which captures both the utility and intra-sample diversity of a sample, and an inter-sample importance score using low-dimensional similarity matching against previously selected samples to quantify inter-sample diversity (D1). Integrating these scores allows the training process to balance exploitation of high-utility samples with exploration of less-visited regions in the data distribution, improving the effectiveness of online batch selection.
@@
    To capture intra-sample value, we leverage the nuclear norm of the logits matrix, which naturally reflects data utility and intra-sample diversity, providing a principled criterion for assessing within-sample informativeness.
@@
    To estimate inter-sample diversity, we design a structured bilinear random projection of logits for compact embedding and use similarity matching with historical data points to reduce redundancy with negligible overhead.
@@
    Our UDS operates directly on forward-pass outputs without external resources, achieving faster convergence than full-dataset SFT and surpassing existing baselines.
@@ 2.1 Problem Formulation of Online Batch Selection
At training iteration $t$, we draw a candidate batch $\mathcal{B}_{t}=\{(\bm{x}_{t}^{i},y_{t}^{i})\}_{i=1}^{B}$ from the corpus. The goal of online batch selection is to choose a subset $\widehat{\mathcal{B}}_{t}\subseteq\mathcal{B}_{t}$ that maximizes the immediate (or near-term) benefit to the model. Let $s(\bm{x}_{t}^{i};\bm{\theta}_{t},\mathcal{H}_{t})\in\mathbb{R}$ denote a per-sample importance score computed under the current model state $\bm{\theta}_{t}$ and optional history $\mathcal{H}_{t}$ (e.g., recent sample embeddings). The selection objective can be viewed as an optimization problem:
@@
where $K\in\{1,\dots,B\}$ is the target batch size.
@@ 2.2 Autoregressive Language Modeling
During fine-tuning, the parameters $\bm{\theta}_{t}$ are updated to maximize the likelihood of training sequences. At iteration $t$, the online selection mechanism identifies the most valuable subset $\widehat{\mathcal{B}}_{t}$ from the candidate batch $\mathcal{B}_{t}$, which is then used to perform a parameter update $\bm{\theta}_{t}\to\bm{\theta}_{t+1}$.
@@ 3 Utility-Diversity Sampling (UDS)
We propose UDS, an online batch selection framework that jointly considers data utility, intra-sample diversity, and inter-sample diversity. It scores and samples each data point using a mixture of two complementary scores: (i) the Nuclear Norm of the logits matrix ($s_{\text{intra}}^{t,i}$), capturing optimization utility and intra-sample diversity; and (ii) the Diversity Distance ($s_{\text{inter}}^{t,i}$), measuring dispersion against historical selections. An overview is shown in Figure 1, with pseudocode provided in Algorithm 1.
@@ 3.1 Intra-sample Importance Score via Nuclear Norm
The intra-sample importance score ($s_{\text{intra}}^{t,i}$) can be characterized through two complementary views: (i) optimization utility—how much the sample can contribute to loss reduction during training, and (ii) intra-sample diversity—how diverse the model’s token-level outputs are within the sequence. Inspired by Cui et al. (2020), we employ the nuclear norm (trace norm) of the logits matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ to capture both aspects. A strong correlation between the nuclear norm $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ and these two aspects is visualized in Figures 2(a) and 2(b). $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$ is defined as the sum of the singular values of $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$. Let $\{\sigma_{1},\sigma_{2},\dots,\sigma_{r}\}$ denote its singular values (with $r=\mathrm{rank}\left(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\right)$), then $s_{\text{intra}}^{t,i}$ is computed as:
@@
In mathematics, the nuclear norm and the Frobenius norm are related via loose bounds, as shown in Lemma 3.1. A detailed explanation is provided in Appendix D.1.
@@
(Horn and Johnson, 2012) For any matrix $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\in\mathbb{R}^{N\times V}$, the following inequality holds:
$\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}\leq\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}\leq C\cdot\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F},$
where $C=\sqrt{\min(N,V)}$ and $\|\cdot\|_{F}$ denotes the Frobenius norm, computed as the square root of the sum of squares of all entries in the matrix. The left inequality holds with equality when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is rank-$1$ (i.e., has only one non-zero singular value). The right inequality holds with equality when $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ has full rank and all singular values are equal.
Due to the significant gap between the lower and upper bounds in Lemma 3.1, a larger nuclear norm can arise in two ways: (1) the Frobenius norm increases, typically indicating larger logits, and (2) for a fixed Frobenius norm, the nuclear norm shifts closer to the upper bound. Below, we intuitively clarify how these two effects connect to the two complementary aspects of the sample-importance score.
@@
Input: Candidate batch sequence $\mathcal{B}_{t}$ for $t=0,\ldots,T$; Initial model $\pi_{\bm{\theta}_{0}}$; Memory queue capacity $M$; Candidate batch size $B$; Target batch size $K$; Trade-off factor $\alpha$.
Output: Finetuned Model $\pi_{\bm{\theta}_{T+1}}$
Initialize memory queue $\bm{Q}\leftarrow\emptyset$
for $t=0$ to $T$ do
 // Calculate Importance Score foreach $\bm{x}_{t}^{i}\in\mathcal{B}_{t}$ do
    $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\leftarrow\pi_{\bm{\theta}_{t}}(\bm{x}_{t}^{i})$$s_{\text{intra}}^{t,i}\leftarrow\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$$\bm{z}_{t}^{i}\leftarrow\operatorname{vec}(\bm{\Gamma}_{2}\cdot\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\cdot\bm{\Gamma}_{1}^{\top})$$s_{\text{inter}}^{t,i}\leftarrow\frac{1}{|\bm{Q}|}\sum_{\bm{z}_{j}\in\bm{Q}}\|\bm{z}_{t}^{i}-\bm{z}_{j}\|_{2}$$s_{\text{total}}^{t,i}\leftarrow s_{\text{intra}}^{t,i}+\alpha\cdot s_{\text{inter}}^{t,i}$
 // Select top-$K$ samples $\widehat{\mathcal{B}}_{t}\leftarrow\operatorname{TopK}(\{s_{\text{total}}^{t,i}\}_{i=1}^{B},K)$// Update Memory and Model while $|\bm{Q}|+K>M$ do
    $\bm{Q}\leftarrow\bm{Q}\setminus\{\text{oldest element}\}$
 $\bm{Q}\leftarrow\bm{Q}\cup\{\bm{z}_{t}^{i}\mid\bm{x}_{t}^{i}\in\widehat{\mathcal{B}}_{t}\}$  Update parameters $\bm{\theta}_{t}$ through backpropagation on $\widehat{\mathcal{B}}_{t}$
Larger Nuclear Norm Correlates With Higher Optimization Utility.
Traditional notions of sample difficulty in online batch selection—such as maximum loss (Loshchilov and Hutter, 2015) or maximum gradient (Katharopoulos and Fleuret, 2018)—often misalign with actual training dynamics (Wang et al., 2024). Instead, we characterize sample importance through its potential contribution to loss reduction, which has been widely used in previous work and which we term optimization utility. For ease of exposition, we assume training with SGD using batch size $B$ and a learning rate $\eta_{t}$. An extension to Adam is provided in Appendix D.2. The parameter update at iteration $t$ is $\bm{\theta}_{t+1}-\bm{\theta}_{t}=-\frac{\eta_{t}}{B}\sum_{j=1}^{B}\nabla_{\bm{\theta}}\ell(\bm{x}_{t}^{j};\bm{\theta}_{t})$. After the update, the logits matrix for datapoint $\bm{x}_{t}^{i}$ can be expanded using a first-order Taylor expansion:
@@
$\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})=\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t+1})-\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$
$\approx-\frac{\eta_{t}}{B}\nabla_{\bm{\theta}}\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\cdot\left(\sum_{j=1}^{B}\nabla_{\bm{\theta}}\ell(\bm{x}_{t}^{j};\bm{\theta}_{t})\right).$ (6)
We can view the training step as producing a logits perturbation $\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$. The first-order Taylor expansion of the loss around $\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ gives
$\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$ $:=\ell\left(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})+\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\right)-\ell\left(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\right)$
$\approx\langle\nabla_{\bm{L}}\ell(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})),\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\rangle,$ (7)
@@
(a) Loss Reduction vs Nuclear Norm
(b) Rank vs Nuclear Norm
(c) Loss Reduction vs Frobenius Norm
Figure 2: Correlation Analysis using Qwen-2.5-7B on MMLU. (a) Strong linear correlation between $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$ and $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$. (b) Strong linear correlation between $\mathrm{rank}(\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t}))$ and $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$. (c) Strong linear correlation between $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$ and $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$. During training optimized with Adam, each point represents the correlation coefficient computed from $B$ pairs per batch.
Due to model complexity and nonlinearity, strictly proving the nuclear norm’s linear correlation with loss reduction is challenging. Instead, we provide an intuitive explanation. According to Equation 8, as the Frobenius norm of the logits $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$ increases, the perturbation $\delta\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ also tends to grow uniformly. This occurs because each entry in the activation logits becomes larger, and these values are directly used during backpropagation. In contrast, $\bm{\Delta}(\bm{x}_{t}^{i};\bm{\theta}_{t})$ is less sensitive to scale because it depends only on the normalized probabilities and the one-hot labels. Therefore, $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{F}$, $\|\bm{L}(\bm{x}_{t}^{i};\bm{\theta}_{t})\|_{*}$, and the attainable loss reduction $\delta\ell(\bm{x}_{t}^{i};\bm{\theta}_{t})$ typically grow in tandem. Empirically, Figure 2(a) and Figure 2(c) confirm a strong correlation between a sample’s loss reduction and its nuclear norm and Frobenius norm, respectively. Hence, the nuclear norm can serve as an indicator of a sample’s optimization utility.
When a single sample induces a larger loss reduction, it indicates that the sample both challenges the model’s current predictions and provides strong gradients that effectively guide parameter updates. Such samples are particularly valuable for data selection: they not only highlight under-learned regions of the input space, but also accelerate training by yielding stronger optimization signals compared to redundant or already well-learned samples. We provide a categorization of the loss before and after training in Table 2. We also provide a discussion from a curriculum learning perspective in Appendix B.2 to analyze what kinds of data are selected as the training evolves. Hence, we prioritize selecting samples with larger nuclear norm that are more likely to have higher loss reduction.
@@
High Too Hard High Utility